Ask
your own question, for FREE!
Mathematics
5 Online
50% of people drive to work, 30% have children, and 20% drive and have children. What is the probability of a person driving and having children the answer is 3/5 but how?
Still Need Help?
Join the QuestionCove community and study together with friends!
=0.3 x 0.2 = 0.6 = 3/5
easy way to think about it. In probability "AND" means multiply and "OR" means you add.
Similar to how logic works.
"..20% drive and have children. What is the probability of a person driving and having children..."
is this a typo or am i not reading it right?
Still Need Help?
Join the QuestionCove community and study together with friends!
thank you amister that is not correct elec that is both and have children
wait
ohhh I think I get it
I think it's conditional probability but the question is not worded correctly.
oh ok srry
Still Need Help?
Join the QuestionCove community and study together with friends!
i am srry it was probability of a person driving OR having children
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
thereneelg:
Okay....I'm doing a group solo for choir. Auditions are Tuesday (?), Any tips on
Arizona:
my Rolling stone airpods won't connect and I got them about a year ago is there t
Arizona:
can someone check out my essay i'm doing in my senior English class about the McD
IcedShots:
I was wondering, is there anybody on here around my age? Im turning 15 on December 1 ud83dude0a.
thatonenerd:
Use Hvap of water = 40.7 kJ/mol, c of water = 4.184 J/g.u00b0C, and molar mass of HO = 18.
letrasbonitasz:
Typography plays an important role in digital communication. Different font styles can make social media profiles, usernames, messages, and creative designs
Bounty:
tested love: im so over love it acquires patience i don't have trials an tribulations yet my souls the only one taken i give 90 you give 10 you say that you
SaltySuga:
is it possible to tell how deep a cut is by an open wound? like would i be able t
AndreW2002:
I'm trying to change my profile picture but it keeps saying "Either a error occurred, or you tried to upload a gif" but it's not a gif that I'm using, it's
10 hours ago
0 Replies
0 Medals
17 hours ago
3 Replies
0 Medals
1 day ago
1 Reply
0 Medals
17 hours ago
3 Replies
0 Medals
2 days ago
5 Replies
0 Medals
4 days ago
3 Replies
1 Medal
2 days ago
7 Replies
2 Medals
5 days ago
5 Replies
1 Medal
6 days ago
3 Replies
0 Medals